OpenStudy (anonymous):

MEDAL What is the equation in point−slope form of the line passing through (2, 2) and (4, 8)

pooja195 (pooja195):

\[\huge\bf \frac{ y_2 - y_1 }{ x_2 -x_1}\] formula to find slope x and y values are \[\large\rm (x_1 ,y_1) (x_2 ,y_2)\]

OpenStudy (anonymous):

so i guess ( y-2)=3(x-2)